Could you please help me with the following question?
Show that if n > 1, then P(n,2)=nē-n

Answer by jim_thompson5910(35256)

P(n,r)=n!/(n-r)! ... Start with the permutation formula


P(n,2)=n!/(n-2)! ... Plug in r=2


P(n,2)=[n(n-1)(n-2)!]/(n-2)! ... Rewrite n! as n(n-1)(n-2)!


P(n,2)=n(n-1) ... Cancel out the common terms (n-2)!


P(n,2)=n^2-n ... Distribute